Keisha's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Keisha
$5.35
per pound, and type B coffee costs
$4.30
per pound. This month, Keisha made
126
pounds of the blend, for a total cost of
$595.35
. How many pounds of type B coffee did she use?

Answer by ewatrrr(24785)

5.35x + 4.30(126-x) = 595.35
x = (595.35-126(4.30)/1.05, amount of A
B = 126-A